4.The cost of building an office complex, x floors high, in a prime location in Accra is made up of three components: 

(a) GH¢10 million for the land

(b) GH¢1/4 million per floor

(c) Specialized costs of GH¢10000x per floor.

How many floors should the office complex contain if the average cost per floor is to be minimized?

5. The supply and demand equations of a good are respectively given by  and 

The government decides to impose a tax, t, per unit. Find the value of t (in Ghana cedis) which maximizes the governments total tax revenue on the assumption that equilibrium conditions prevail in the market. 

6. A firm's demand function for a certain good is given by. Its total cost function is . What output level maximizes the firm's profit? 


1
Expert's answer
2021-05-16T17:42:20-0400

4. Average cost per floor ATC is minimized if ATC = MC.

ATC = TC/x = (10 + 0.26x)/x = 10/x + 0.26.

MC = TC'(x) = 0.26.

10/x + 0.26 = 0.26,

10/x = 0.

There is no root, so, there is no point, at which ATC is minimized.

5. Total tax revenue is maximized when marginal tax revenue equals zero.

6. The firm's profit is maximized at such output level, for which MR = MC.